What is the amount of charge possessed by 1kv of electrons?Given that mass of an electronelis 91x10-31kg
 

Grade:12

1 Answers

Arun
4 years ago
Charge of an electron: -1.602 * 10^-19 C
Mass of an electron: 9.11 * 10^-31 kg
Electrons in 1kg: 1/(9.11 * 10^-31) = 1.10 * 10^30.
Charge of 1kg of electrons: (1.10 * 10^30)(-1.602 * 10^-19) = -1.76 * 10^11 C